Prove that

x<log(1/1-x)< x/1-x, 0<x<1



1
Expert's answer
2021-06-15T03:40:28-0400

Solution:

Let "f(x)=\log (\frac1{1-x})" in [0, 1]

Since f(x) satisfies the condition of L.M.V. theorem in [0, 1], there exists "\theta(0<\theta<1)" such that

"\dfrac{f(x)-f(0)}{x-0}=f^{\prime}(\theta x)
\\
\Rightarrow \quad \dfrac{\log (\frac1{1-x})}{x}=\dfrac{}{}\dfrac{1}{1-\theta x}"

"\Rightarrow \quad \log (\frac1{1-x})=\dfrac{x}{1-\theta x} \quad \ldots (i)"

Now, "\quad 0<\theta<1, 0<x<1 
\Rightarrow \theta x<x"

"\Rightarrow \quad -\theta x>-x
\\\Rightarrow \quad 1-\theta x>1-x
\\
\Rightarrow \quad \frac{1}{1-\theta x}<\frac{1}{1-x}
\\\Rightarrow \quad \frac{x}{1-\theta x}<\frac{x}{1-x}
 \quad \ldots (ii)"

Again "\quad 0<\theta<1, 0<x<1"

"\Rightarrow \quad \theta x>0
\\\Rightarrow \quad -\theta x<0
\\ \Rightarrow \quad 1-\theta x<1
\\
\Rightarrow \quad \frac{1}{1-\theta x}>1
\\
\Rightarrow \quad \frac{x}{1-\theta x}>x \quad \ldots (iii)"

From (i), (ii) and (iii), we get, "x< \log (\frac1{1-x})<\dfrac{x}{1-x}, 0<x<1"